# Accelerated Gradient Methods with Biased Gradient Estimates: Risk Sensitivity, High-Probability Guarantees, and Large Deviation Bounds

### v1

@@
Abstract. We study trade-offs between convergence rate and robustness to gradient errors in the context of first-order methods. Our focus is on generalized momentum methods (GMMs)—a broad class that includes Nesterov’s accelerated gradient, heavy-ball, and gradient descent methods— for minimizing smooth strongly convex objectives. We allow stochastic gradient errors that may be adversarial and biased, and quantify robustness of these methods to gradient errors via the risk-sensitive index (RSI) from robust control theory. For quadratic objectives with i.i.d. Gaussian noise, we give closed form expressions for RSI in terms of solutions to $2\times$ 2 matrix Riccati equations, revealing a Pareto frontier between RSI and convergence rate over the choice of stepsize and momentum parameters. We then prove a large-deviation principle for time-averaged suboptimality in the large iteration limit and show that the rate function is, up to a scaling, the convex conjugate of the RSI function. We further show that the rate function and RSI are linked to the $H_{\infty}$-norm—a measure of robustness to the worst-case deterministic gradient errors—so that stronger worst-case robustness (smaller $H_{\infty}$ norm) leads to sharper decay of the tail probabilities for the average suboptimality. Beyond quadratics, under potentially biased sub-Gaussian gradient errors, we derive non-asymptotic bounds on a finite-time analogue of the RSI, yielding finite-time high-probability guarantees and non-asymptotic large-deviation bounds for the averaged iterates. In this more general scenario of smooth strongly convex functions, we also observe an analogous trade-off between RSI and convergence-rate bounds. To our knowledge, these are the first non-asymptotic guarantees for GMMs with biased gradients and the first risk-sensitive analysis of GMMs. Finally, we provide numerical experiments on a robust regression problem to illustrate our results.
@@
footnotetext: The authors can be contacted at the following email addresses: mg1366@rutgers.edu (M. Gürbüzbalaban), 🖂 yasa.syed@rutgers.edu (Y. Syed), 🖂 nsa10@psu.edu (N. S. Aybat).
@@ 1. Introduction
When gradient errors are persistent or decay slowly, the iterates may fail to converge—oscillating around the optimal solution, deviating significantly before eventually converging, or even diverging entirely [16, 17, 14, 49]. Robustness as the sensitivity to gradient errors, measured in terms of some properly defined solution accuracy metric, is a critical performance criterion [14, 50]. In particular, while momentum-based methods such as NAG, HB, and TMM are known to converge faster than the standard GD in noise-free settings for convex or strongly convex problems, they are generally less robust to gradient errors [14, 13, 49, 51]. As a result, designing effective algorithm parameters requires a careful understanding of the trade-offs between convergence speed and robustness.
Existing robustness analyses of momentum methods have primarily concentrated on two extremes. On one side, a growing line of work investigates stochastic, unbiased noise, establishing mean-square or high-probability convergence guarantees under light-tailed assumptions (e.g., [52, 53, 54, 55, 56]). On the other side, control-theoretic approaches employ the $H_{\infty}$-norm to characterize worst-case deterministic gradient errors [57], thereby modeling the gradient as deterministically biased. However, $H_{\infty}$-based analyses usually ignore stochastic variability and so far lack high-probability guarantees on the iteration complexity for obtaining approximate solutions under stochastically biased gradients.
@@
$R_{K}(\theta)=\frac{2\sigma^{2}}{\theta\,(K+1)}\log\left(\mathbb{E}\exp\left[\frac{\theta}{2\sigma^{2}}\,S_{K}\right]\right),\quad S_{K}=\sum_{k=0}^{K}\left(f(x_{k})-f_{*}\right),$ (1.1)
@@
The parameter $\theta>0$ in the cost $R_{K}(\theta)$ provides a continuous interpolation between the risk-neutral regime—corresponding to the limit $\theta\to 0$, capturing the average performance $\mathbb{E}\!\left[\tfrac{S_{K}}{K+1}\right]$—and progressively more risk-averse regimes that place increasing emphasis on tail events for larger positive values of $\theta$ and for sufficiently large $\theta$, approximating worst-case behavior [62]. Conversely, $\theta<0$ induces a risk-seeking regime that down-weights large suboptimality events and favors variability. Thus, by varying $\theta$, the risk-sensitive cost spans average-case, risk-seeking, and worst-case perspectives with a unified lens. In addition, the risk-sensitive cost $R_{K}(\theta)$ is a scaled cumulant-generating function, i.e., natural logarithm of the moment generating function, and Chernoff bounds turn $R_{K}(\theta)$ into explicit exponential tails for sub-optimality, yielding high-probability bounds and rate functions that characterize the decay rate for the average suboptimality $\mathbb{P}\!\left[\tfrac{S_{K}}{K+1}>t\right]$. These properties of the risk-sensitive index make it interesting to study in the context of optimization algorithms.
@@
To the best of our knowledge, this work provides the first risk-sensitive theory and analysis of generalized momentum methods. Unlike prior research, which has been predominantly risk-neutral and has typically relied on the assumption of unbiased stochastic gradients, our framework accommodates gradient noise that may be biased, thereby delivering high-probability guarantees that remain valid under biased gradient settings.
@@
Contributions. Our main contributions are built on several pillars. First, we analyze the strongly convex quadratic setting under i.i.d. isotropic Gaussian gradient noise. This case provides a clean and tractable foundation, allowing us to develop explicit characterizations before addressing the more general class of smooth, strongly convex functions subject to potentially biased noise. In the quadratic setting, we investigate how the risk-sensitive index $R(\theta)$ depends on the given algorithm parameters $(\alpha,\beta,\nu)$ and risk-averseness parameter $\theta$. By modeling the algorithmic class GMM as a linear time-invariant dynamical system and leveraging some existing results from robust control theory, we argue that $R(\theta)$ admits an explicit characterization through solving a discrete-time algebraic Riccati equation (DARE) that depends on the algorithm parameters. However, this is a matrix equation over $2d\times 2d$ matrices, and in general, it entails solving a $4d\times 4d$ generalized eigenvalue problem, which is computationally prohibitive for large $d$ requiring at least $\mathcal{O}(d^{3})$ operations. To this end, we establish a dimension-reduction result: for quadratic objectives subject to i.i.d. isotropic Gaussian noise with variance $\sigma^{2}$, the DARE decomposes into $d$ independent $2\times 2$ Riccati equations. Each such equation can be solved in $\mathcal{O}(1)$ time, allowing the overall solution to be computed in $\mathcal{O}(d)$ operations in parallel. This yields a closed-form expression for the risk-sensitive index in terms of easily solvable $2\times 2$ DAREs. Furthermore, for gradient descent, we can solve the DAREs explicitly and provide an explicit expression for the risk-sensitive index. Building on our characterization of the risk-sensitive index, we then establish a fundamental trade-off between risk (as measured by this index) and convergence rate (quantifying solution speed in the absence of noise) in momentum methods. In particular, parameter choices that slow down convergence speed also reduce the associated risk and enlarge the range of $\theta$ values for which the risk remains finite, i.e., improved risk sensitivity and greater flexibility in $\theta$ comes at the expense of a reduction in the convergence rate. Specifically, the risk is finite if and only if $\theta>0$ satisfies $\sqrt{\theta}\,H_{\infty}<\sqrt{d},$ where $H_{\infty}$ denotes the $H_{\infty}$-norm of the dynamical system associated with the GMM iterations—explicitly characterized for quadratic objectives in [57]. Because the $H_{\infty}$-norm quantifies robustness to the worst-case deterministic gradient errors [57], this relation implies that greater robustness in the $H_{\infty}$-sense permits a wider range of $\theta$ for which the risk is finite, thereby linking risk sensitivity (a probabilistic performance measure) to deterministic worst-case robustness in momentum-based methods— to the best of our knowledge, for the first time. While connections between risk sensitivity, and the $H_{\infty}$ norm exists in the context of control systems [61, 63], what these connections mean in the context of momentum-based optimization algorithms subject to gradient errors are first explored in our work to our knowledge. We then compare momentum methods—HB, NAG, GD, and TMM—across different parameter settings, focusing on their risk profiles, i.e., how their risk sensitivity varies with stepsize and momentum. We find that HB with standard parameters (proposed by Polyak [8]) converges faster than NAG (with standard parameters as in [7]) but incurs higher risk, whereas GD (with the common stepsize $\alpha=1/L$ where $L$ is the Lipschitz constant of the gradient) is slower yet exhibits a lower risk than both methods. These trade-offs trace a Pareto-optimal frontier over parameter choices, delivering the best attainable risk for a target convergence rate –we illustrate this frontier numerically by leveraging our dimension-reduction framework to compute the risk index.
@@
Second, we establish a large-deviation principle for the average suboptimality $\Delta_{K}\coloneqq S_{K}/(K+1)$, quantifying the exponential decay rate of rare-event probabilities for $\Delta_{K}$ in the quadratic setting. This also allows us to control the probability of rare events associated with suboptimality of the ergodic averages of the iterates. We show that the rate function is the Fenchel–Legendre transform of a scaled version of the risk index $R(\theta)$. A key challenge to establish such a result is that the summands $f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}$ that appear in the sum $S_{K}$ are neither independent nor identically distributed—even when the gradient noise $\{w_{k+1}\}_{k\geq 0}$ is i.i.d.)—so classical results such as Cramér’s theorem do not apply. Instead, we verify the hypotheses of the Gärtner–Ellis theorem [64], which accommodates such non-i.i.d. sequences, to identify the rate function. The analysis takes advantage of our characterizations of the risk-sensitive index for momentum methods. These results demonstrate that the rate function and the risk index—which govern the tail behavior of suboptimality in the large-iteration limit—are intrinsically linked to the $H_{\infty}$ norm. Moreover, they show that greater robustness to worst-case deterministic errors (as quantified by the $H_{\infty}$ norm) is associated with a sharper decay of tail probabilities, reflected in a larger rate function.
@@
In our third set of results, we extend the analysis to the minimization of strongly convex and smooth. In this setting, explicit characterizations of the risk index and the $H_{\infty}$ norm are generally unavailable. While upper bounds $\overline{H}_{\infty}$ exist for the $H_{\infty}$ norm [57], no analogous bounds are known for the risk-sensitive cost or the risk index. To bridge this gap, we consider a more general noise model in which the gradient errors are sub-Gaussian, and potentially biased. This model extends the Gaussian noise considered in the quadratic setting and accommodates stochastic errors with sub-Gaussian tails, including bounded stochastic errors, deterministic errors, and their combinations. We derive upper bounds on $R_{K}(\theta)$ at every iteration $K$, and translate these into finite-time high-probability guarantees and non-asymptotic large-deviation upper bounds for the averaged iterates. Our approach relies on introducing a novel Lyapunov function together with a backward-induction argument, inspired by dynamic programming methods, which integrates the noise to iteratively estimate the risk-sensitive cost. Taking the limit as $K\to\infty$, we derive bounds on the risk index $R(\theta)$ and associated large-deviation upper bounds. These results can be interpreted as non-asymptotic counterparts of the asymptotic characterizations established in the quadratic case. To the best of our knowledge, such non-asymptotic high-probability and large-deviation bounds for biased gradient estimates in the general class of momentum methods considered here have not previously appeared in the literature. Analogous to the quadratic case, we establish intrinsic trade-offs between risk and convergence rate in the design of first-order accelerated methods. Moreover, we find that the bounds on the $H_{\infty}$ norm controls the range of $\theta$ values for which the risk remains finite. Our bounds connect $R(\theta)$, high-probability guarantees, and large-deviation rates to the $H_{\infty}$ norm, thereby sharpening the link between probabilistic performance and worst-case robustness. They show that greater robustness to deterministic errors (as quantified by the $H_{\infty}$ norm) is associated with sharper tail decay in average suboptimality and reduced bias in the averaged iterates.
@@ 2. Related Work
Under a light-tailed (sub-Gaussian) noise assumption, high-probability convergence guarantees for stochastic gradient descent (SGD) and its momentum variants have been established for convex objectives [24, 65, 66, 67, 68, 69, 70, 71] and for non-convex objectives [72, 73, 67]. Beyond this setting, several works relax the light-tailed assumption and instead provide guarantees under bounded variance noise [74, 75] or, more generally, sub-Weibull noise [76]. Unlike our setting, however, these works assume unbiased gradient noise and do not extend to the general class of momentum methods (GMM) studied in this paper.
@@
Compared to in-expectation guarantees, high-probability guarantees under biased noise are less developed. While in-expectation guarantees can be converted into high-probability guarantees using strategies such as longer runs of the algorithm combined with Chebyshev’s inequality, or parallel runs with selection of the best outcome [79, p. 243],[80], these strategies often produce loose complexity bounds with respect to the target probability and desired precision and can be impractical in streaming data settings [81],[56, Remark 13]. Existing high-probability results include random reshuffling for SGD [31, 30] and non-linear SGD variants with heavy-tailed noise where clipping, quantization, or normalization induce bias [82, 67].
@@
Next, we are going to discuss a related body of work from the risk perspective. Risk sensitivity, and more broadly, entropic risk, has been covered in the literature largely in the control theory context [59, 91, 60, 61]. These works consider centered, independent identically distributed Gaussian noise. In particular, in [61], an exact analysis of $R(\theta)$ is provided in the linear–quadratic control setting. These results are not given in the context of optimization algorithms but rather for general linear dynamical systems. In our work, however, we interpret GMM as a linear dynamical system when the objective is a quadratic and show that their result translates directly into a formula for the risk-sensitive cost of GMM algorithms in this setting. However, computing the risk-sensitivity index via the approach of [61] requires integrating the system’s transfer function in the frequency domain, which—using the results of [63], [92, Remark 4.11], and [93, Lemma 3.4]—can be reduced to solving $2d\times 2d$ Riccati equations, entailing at least $\mathcal{O}(d^{3})$ computational complexity. By exploiting the special structure of the GMM, we show that it suffices to solve $2\times 2$ Riccati equations, which involve three unknowns—the entries of a symmetric $2\times 2$ matrix—and three corresponding equations; moreover, the approach in [61] applies only to quadratics and does not extend to strongly convex objectives. More recently, in [53] the authors introduced robustness measures related to asymptotic optimality under unbiased gradient errors, analyzing quadratic functions and, more generally, smooth strongly convex functions. In particular, the trade-offs between convergence rate and robustness to gradient errors for GD and NAG has been analyzed in [53]; however, in contrast to ours, the approach in [53] does not extend to biased gradient estimates or deterministic worst-case errors –the rate vs risk trade-off scheme introduced in [53] for first-order optimization methods is later extended in [55] to analyze first-order primal-dual methods for solving strongly convex-strongly concave saddle point problems subject to unbiased gradient errors with finite variance. In a different direction from [53], a more recent work [58] proposes an alternative robustness measure motivated by the entropic risk of the iterates and provide high-probability bounds for GMM in the unbiased sub-Gaussian gradient noise setting. However, their analysis does not accommodate biased gradient estimates, focuses solely on the suboptimality of the last iterate $x_{K}$ rather than on ergodic averages or cumulative suboptimality $S_{K}$, and does not establish large-deviation bounds.
@@
$\frac{\mu}{2}\|x-y\|^{2}\leq f(x)-f(y)-\nabla f(y)^{\top}(y-x)\leq\frac{L}{2}\|x-y\|^{2},\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\forall~x,y\in\mathbb{R}^{d}.}$ (2.1)
Due to strong convexity, $f$ admits a unique global minimum on $\mathbb{R}^{d}$ which we will denote by $x_{*}$. From (2.1), one has $\mu\leq L$, and we assume $\mu<L$ throughout the paper –otherwise, for $\mu=L$, the class $\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$ is trivial, i.e., $f(x)=\frac{\mu}{2}\|x\|^{2}$. Let $\mathsf{I}_{d}$ denote the $d\times d$ identity and let $0_{d}$ be the $d\times d$ zero matrix. We drop the subscript $d$ in some cases when it is clear from the context. Given matrix $A$, let $A_{ij}$ denote the entry on the $i$-th row and $j$-th column of $A$. The spectral radius $\rho(A)$ of a square matrix $A$ is defined as the largest modulus of the eigenvalues of $A$, and $\|A\|$ denotes the spectral norm of $A$. Throughout the paper, $A\otimes B$ denotes the Kronecker product of the matrices $A$ and $B$.

### v2

# and Large Deviation Bounds

@@
Abstract. We study trade-offs between convergence rate and robustness to gradient errors in the context of first-order methods. Our focus is on generalized momentum methods (GMMs)—a broad class that includes Nesterov’s accelerated gradient, heavy-ball, and gradient descent methods— for minimizing smooth strongly convex objectives. We allow stochastic gradient errors that may be adversarial and biased, and quantify robustness of these methods to gradient errors via the risk-sensitive index (RSI) from robust control theory. For quadratic objectives with i.i.d. Gaussian noise, we give closed form expressions for RSI in terms of solutions to $2\times$ 2 matrix Riccati equations, revealing a Pareto frontier between RSI and convergence rate over the choice of stepsize and momentum parameters. We then prove a large-deviation principle for time-averaged suboptimality in the large iteration limit and show that the rate function is, up to a scaling, the convex conjugate of the RSI function. We further show that the rate function and RSI are linked to the $H_{\infty}$-norm—a measure of robustness to the worst-case deterministic gradient errors—so that stronger worst-case robustness (smaller $H_{\infty}$-norm) leads to sharper decay of the tail probabilities for the average suboptimality. Beyond quadratics, under potentially biased sub-Gaussian gradient errors, we derive non-asymptotic bounds on a finite-time analogue of the RSI, yielding finite-time high-probability guarantees and non-asymptotic large-deviation bounds for the averaged iterates. In this more general scenario of smooth strongly convex functions, we also observe an analogous trade-off between RSI and convergence-rate bounds. To our knowledge, these are the first non-asymptotic guarantees for GMMs with biased gradients and the first risk-sensitive analysis of GMMs. Finally, we provide numerical experiments on a robust regression problem to illustrate our results.
@@ 1. Introduction
When gradient errors are persistent or decay slowly, the iterates may fail to converge—oscillating around the optimal solution, deviating significantly before eventually converging, or even diverging entirely [16, 17, 14, 49]. Robustness—understood as sensitivity to gradient errors and quantified by an appropriate solution-accuracy metric—is therefore a critical performance criterion [14, 50]. In particular, while momentum-based methods such as NAG, HB, and TMM are known to converge faster than the standard GD in noise-free settings for convex or strongly convex problems, they are generally less robust to gradient errors [14, 13, 49, 51]. As a result, designing effective algorithm parameters requires a careful understanding of the trade-offs between convergence speed and robustness.
Existing robustness analyses of momentum methods have primarily concentrated on two extremes. On one end, a growing line of work investigates stochastic, unbiased noise, establishing mean-square or high-probability convergence guarantees under light-tailed assumptions (e.g., [52, 53, 54, 55, 56]). On the other end, control-theoretic approaches employ the $H_{\infty}$-norm to characterize worst-case deterministic gradient errors [57], thereby modeling the gradient as deterministically biased. However, $H_{\infty}$-based analyses usually ignore stochastic variability and typically do not provide high-probability guarantees on the iteration complexity for obtaining approximate solutions under stochastically biased gradients.
@@
$R_{K}(\theta)=\frac{2\sigma^{2}}{\theta\,(K+1)}\log\left(\mathbb{E}\exp\left[\frac{\theta}{2\sigma^{2}}\,S_{K}\right]\right),\qquad S_{K}=\sum_{k=0}^{K}\left(f(x_{k})-f_{*}\right),$ (1.1)
@@
The parameter $\theta>0$ in the cost $R_{K}(\theta)$ provides a continuous interpolation between the risk-neutral regime—corresponding to the limit $\theta\to 0$, capturing the average performance $\mathbb{E}\!\left[\tfrac{S_{K}}{K+1}\right]$—and progressively more risk-averse regimes that place increasing emphasis on tail events for larger positive values of $\theta$, and approximating the worst-case behavior for sufficiently large $\theta$ [62]. Conversely, $\theta<0$ induces a risk-seeking regime that down-weights large suboptimality events and favors variability. Thus, by varying $\theta$, the risk-sensitive cost spans average-case, risk-seeking, and worst-case perspectives with a unified lens. In addition, the risk-sensitive cost $R_{K}(\theta)$ is a scaled cumulant-generating function, i.e., natural logarithm of the moment generating function, and Chernoff bounds turn $R_{K}(\theta)$ into explicit exponential tails for suboptimality, yielding high-probability bounds and rate functions that characterize the decay rate for the average suboptimality $\mathbb{P}\!\left[\tfrac{S_{K}}{K+1}>t\right]$. These properties of the risk-sensitive index make it interesting to study in the context of optimization algorithms.
@@
To the best of our knowledge, this work provides the first risk-sensitive analysis of generalized momentum methods. Unlike prior research, which has been predominantly risk-neutral and has typically relied on the assumption of unbiased stochastic gradients, our framework accommodates gradient noise that may be biased, thereby delivering high-probability guarantees that remain valid under biased gradient settings.
@@
Contributions. Our main contributions are built on several pillars. First, we analyze the strongly convex quadratic setting under i.i.d. isotropic Gaussian gradient noise. This case provides a clean and tractable foundation, allowing us to develop explicit characterizations before addressing the more general class of smooth, strongly convex functions subject to potentially biased noise. In the quadratic setting, we investigate how the risk-sensitive index $R(\theta)$ depends on the given algorithm parameters $(\alpha,\beta,\nu)$ and risk-aversion parameter $\theta$. By modeling the algorithmic class GMM as a linear time-invariant dynamical system and leveraging some existing results from robust control theory, we argue that $R(\theta)$ admits an explicit characterization through solving a discrete-time algebraic Riccati equation (DARE) that depends on the algorithm parameters. However, this is a matrix equation over $2d\times 2d$ matrices, and in general, it entails solving a $4d\times 4d$ generalized eigenvalue problem, which is computationally prohibitive for large $d$ requiring at least $\mathcal{O}(d^{3})$ operations. To this end, we establish a dimension-reduction result: for quadratic objectives subject to i.i.d. isotropic Gaussian noise with variance $\sigma^{2}$, we show that the DARE decomposes into $d$ independent $2\times 2$ Riccati equations. Each such equation can be solved in $\mathcal{O}(1)$ time, allowing the overall solution to be computed in $\mathcal{O}(d)$ operations in parallel. This yields a closed-form expression for the risk-sensitive index in terms of easily solvable $2\times 2$ DAREs. Furthermore, for gradient descent, we can solve the DAREs explicitly and provide an explicit expression for the risk-sensitive index. Building on our characterization of the risk-sensitive index, we then establish a fundamental trade-off between risk (as measured by this index) and convergence rate (quantifying solution speed in the absence of noise) in momentum methods. In particular, parameter choices that slow down convergence speed also reduce the associated risk and enlarge the range of $\theta$ values for which the risk remains finite, i.e., improved risk sensitivity and greater flexibility in $\theta$ comes at the expense of a reduction in the convergence rate. Specifically, the risk is finite if and only if $\theta>0$ satisfies $\sqrt{\theta}\,H_{\infty}<\sqrt{d},$ where $H_{\infty}$ denotes the $H_{\infty}$-norm of the dynamical system associated with the GMM iterations—explicitly characterized for quadratic objectives in [57]. Because the $H_{\infty}$-norm quantifies robustness to the worst-case deterministic gradient errors [57], this relation implies that greater robustness, i.e., smaller $H_{\infty}$-norm, permits a wider range of $\theta$ for which the risk is finite, thereby linking risk sensitivity (a probabilistic performance measure) to deterministic worst-case robustness in momentum-based methods. While connections between risk sensitivity and the $H_{\infty}$-norm exist in the context of control systems and controlled Markov processes [63, Sec. VI], [61, 64]; what these connections mean in the context of momentum-based optimization algorithms subject to gradient errors is first explored in our work to our knowledge. We then compare momentum methods—HB, NAG, GD, and TMM—across different parameter settings, focusing on their risk profiles, i.e., how their risk sensitivity varies with stepsize and momentum. We find that HB with standard parameters (proposed by Polyak [8]) converges faster than NAG (with standard parameters as in [7]) but incurs higher risk, whereas GD (with the common stepsize $\alpha=1/L$ where $L$ is the Lipschitz constant of the gradient) is slower yet exhibits a lower risk than both methods. These trade-offs trace a Pareto-optimal frontier over parameter choices, delivering the best attainable risk for a target convergence rate –we illustrate this frontier numerically by leveraging our dimension-reduction framework to compute the risk index.
@@
Second, we establish a large-deviation principle for the average suboptimality $\Delta_{K}\coloneqq S_{K}/(K+1)$, quantifying the exponential decay rate of rare-event probabilities for $\Delta_{K}$ in the quadratic setting. This also allows us to control the probability of rare events associated with suboptimality of the ergodic averages of the iterates. We show that the rate function is the Fenchel–Legendre transform of a scaled version of the risk-sensitive index $R(\theta)$. A key challenge to establish such a result is that the summands $f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}$ that appear in the sum $S_{K}$ are neither independent nor identically distributed—even when the gradient noise $\{w_{k+1}\}_{k\geq 0}$ is i.i.d.)—so classical results such as Cramér’s theorem do not apply. Instead, we verify the hypotheses of the Gärtner–Ellis theorem [65], which accommodates such non-i.i.d. sequences, to identify the rate function. The analysis takes advantage of our characterizations of the risk-sensitive index for momentum methods. These results demonstrate that the rate function and the risk-sensitive index—which govern the tail behavior of suboptimality in the large-iteration limit—are intrinsically linked to the $H_{\infty}$-norm. Moreover, they show that greater robustness to the worst-case deterministic errors (as quantified by the $H_{\infty}$-norm) is associated with a sharper decay of tail probabilities, reflected in a larger rate function.
@@
In our third set of results, we extend the analysis to the minimization of strongly convex and smooth functions. In this setting, explicit characterizations of the risk-sensitive index and the $H_{\infty}$-norm are generally unavailable. While upper bounds $\overline{H}_{\infty}$ exist for the $H_{\infty}$-norm [57], no analogous bounds are known for the risk-sensitive cost or the risk-sensitive index. To bridge this gap, we adopt a more general noise model in which gradient errors are sub-Gaussian and may be biased. This model subsumes the Gaussian noise studied in the quadratic case and accommodates stochastic errors with sub-Gaussian tails—including bounded stochastic errors, deterministic errors, and their mixtures. We derive upper bounds on $R_{K}(\theta)$ at every iteration $K$, and translate these into finite-time high-probability guarantees and non-asymptotic large-deviation upper bounds for the averaged iterates. Our approach relies on introducing a novel Lyapunov function together with a backward-induction argument, inspired by dynamic programming methods, to iteratively estimate the risk-sensitive cost. Taking the limit as $K\to\infty$, we derive bounds on the risk-sensitive index $R(\theta)$ and associated large-deviation upper bounds. These results can be interpreted as non-asymptotic counterparts of the asymptotic characterizations established in the quadratic case. To the best of our knowledge, such non-asymptotic high-probability and large-deviation bounds for biased gradient estimates in the general class of momentum methods considered here have not previously appeared in the literature. Analogous to the quadratic case, we establish intrinsic trade-offs between risk and convergence rate in the design of first-order accelerated methods. Moreover, we find that the bounds on the $H_{\infty}$ norm control the range of $\theta$ values for which the risk remains finite. Our bounds connect $R(\theta)$, high-probability guarantees, and large-deviation rates to the $H_{\infty}$-norm, thereby sharpening the link between probabilistic performance and worst-case robustness beyond the quadratic setting as well. They show that greater robustness to deterministic errors (as quantified by the $H_{\infty}$-norm) is associated with sharper tail decay in average suboptimality and reduced bias in the averaged iterates.
@@ 2. Related Work
Under a light-tailed (sub-Gaussian) noise assumption, high-probability convergence guarantees for stochastic gradient descent (SGD) and its momentum variants have been established for convex objectives [24, 66, 67, 68, 69, 70, 71, 72] and for non-convex objectives [73, 74, 68]. Beyond this setting, several works relax the light-tailed assumption and instead provide guarantees under bounded variance noise [75, 76] or, for sub-Weibull noise [77]. Unlike our setting, however, these works assume unbiased gradient noise and do not extend to the general class of momentum methods (GMM) studied in this paper.
@@
Compared to in-expectation guarantees, high-probability guarantees under biased noise are less developed. While in-expectation guarantees can be converted into high-probability guarantees using strategies such as longer runs of the algorithm combined with Chebyshev’s inequality, or parallel runs with selection of the best outcome [80, p. 243], [81], these strategies often produce loose complexity bounds with respect to the target probability and desired precision and can be impractical in streaming data settings [82], [56, Remark 13]. Existing high-probability results include random reshuffling for SGD [31, 30] and non-linear SGD variants with heavy-tailed noise where clipping, quantization, or normalization induce bias [83, 68].
@@
Next, we are going to discuss a related body of work from the risk perspective. Risk sensitivity, and more broadly, entropic risk, has been covered in the literature largely in the control theory context [59, 92, 60, 61]. These works consider centered, independent and identically distributed (i.i.d.) Gaussian noise. In particular, in [61], an exact analysis of $R(\theta)$ is provided in the linear–quadratic control setting. These results are not given in the context of optimization algorithms but rather for general linear dynamical systems. In our work, however, we interpret GMM as a linear dynamical system when the objective is a quadratic and show that their result translates directly into a formula for the risk-sensitive cost of GMM algorithms in this setting. However, computing the risk-sensitivity index via the approach of [61] requires integrating the system’s transfer function in the frequency domain, which—using the results of [64], [93, Remark 4.11], and [94, Lemma 3.4]—can be reduced to solving $2d\times 2d$ Riccati equations, entailing at least $\mathcal{O}(d^{3})$ computational complexity. By exploiting the special structure of the GMM, we show that it suffices to solve $2\times 2$ Riccati equations, which involve three unknowns—the entries of a symmetric $2\times 2$ matrix—and three corresponding equations; moreover, the approach in [61] applies only to quadratics and does not extend to strongly convex objectives. More recently, in [53] the authors introduced robustness measures related to asymptotic optimality under unbiased gradient errors, analyzing quadratic functions and, more generally, smooth strongly convex functions. In particular, the trade-offs between convergence rate and robustness to gradient errors for GD and NAG have been analyzed in [53]; however, in contrast to ours, the approach in [53] does not extend to biased gradient estimates or deterministic worst-case errors –the rate vs risk trade-off scheme introduced in [53] for first-order optimization methods is later extended in [55] to analyze first-order primal-dual methods for solving strongly convex-strongly concave saddle point problems subject to unbiased gradient errors with finite variance. In a different direction from [53], a more recent work [58] proposes an alternative robustness measure motivated by the entropic risk of the iterates and provides high-probability bounds for GMM in the unbiased sub-Gaussian gradient noise setting. However, their analysis does not accommodate biased gradient estimates, focuses solely on the suboptimality of the last iterate $x_{K}$ rather than on ergodic averages or cumulative suboptimality $S_{K}$, and does not establish large-deviation bounds. Compared to prior work, we treat (i) constant step sizes (not decaying), (ii) biased sub-Gaussian gradient errors (not just unbiased), and (iii) risk-sensitive metrics (RSI/LDP) and their link to $H_{\infty}$-norm in the context of momentum-based optimization algorithms.
@@
$\frac{\mu}{2}\|x-y\|^{2}\leq f(x)-f(y)-\nabla f(y)^{\top}(y-x)\leq\frac{L}{2}\|x-y\|^{2},\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\forall~x,y\in\mathbb{R}^{d}.}$ (2.1)
Due to strong convexity, $f$ admits a unique global minimum on $\mathbb{R}^{d}$ which we will denote by $x_{*}$. From (2.1), one has $\mu\leq L$, and we assume $\mu<L$ throughout the paper –otherwise, for $\mu=L$, the class $\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$ is trivial, i.e., $f(x)=\frac{\mu}{2}\|x\|^{2}$. Let $\mathsf{I}_{d}$ denote the $d\times d$ identity and let $0_{d}$ be the $d\times d$ zero matrix. We drop the subscript $d$ in some cases when it is clear from the context. Given matrix $A$, let $A_{ij}$ denote the entry on the $i$-th row and $j$-th column of $A$. The spectral radius $\rho(A)$ of a square matrix $A$ is defined as the largest modulus of the eigenvalues of $A$, and $\|A\|$ denotes the spectral norm of $A$. We use $\det(A)$ and $\operatorname{trace}(A)$ to denote the determinant and the trace of a matrix $A$. Throughout the paper, $A\otimes B$ denotes the Kronecker product of the matrices $A$ and $B$.